OpenStudy (anonymous):

After how many months will Rita pay $200 in simple annual interest on a credit card that charges 12% on her $400 balance?

ganeshie8 (ganeshie8):

200 = P*R*T/12 200 = 400 * .12 * T/12 200 = 48 * .1 * T 200/4.8 = T